# Dynamics of defects and interfaces for interacting quantum hard disks

### v1

@@ Acknowledgments
V.D.N. and F.B.T. contributed equally to this work. We thank Juan Garrahan, Tom Ben Ami, and Roderich Moessner for insightful discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 853443). This work was supported by the German Research Foundation DFG via project 499180199 (FOR 5522) and via project 492547816 (TRR 360).
@@
The data to generate all figures in this article is available in Zenodo [Quantum_hard_disks_dataset].
@@ Appendix A Classical Dynamics of Free Quantum Hard Disks
Appendix B Quantum Many-Body Scars Across the Entire Hilbert Space
Fig. 5 illustrates the entanglement entropy as a function of eigenenergy density, with data points colored by the normalized EA order parameter. This visualization specifically focuses on the largest fragment within the weak fragmentation regime. To extend this analysis, Fig. 7 presents the same plot of entanglement entropy versus eigenenergy density for all fragments. This reveals that numerous additional scar states emerge from the smaller fragments, complementing those found in the largest fragment. Notably, the scars observed in the largest fragment primarily stem from quantum interference effects, whereas those in the smaller fragments emerge due to the fragmentation structure itself.
@@ Appendix B Quantum Many-Body Scars Across the Entire Hilbert Space
Figure 7: Quantum Many-Body Scars across the entire Hilbert space. The entanglement entropy of the eigenstates of all fragments is shown for different interaction strengths: (A) $\lambda=0$, representing free hard disks; (B) $\lambda=0.01$, showing weak interactions; (C) $\lambda=0.1$, illustrating moderate interactions; and (D) $\lambda=0.8$, highlighting strong interaction effects.

### v2

@@ Acknowledgments
We thank Juan Garrahan, Tom Ben-Ami, and Roderich Moessner for insightful discussions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 853443). This work was supported by the German Research Foundation DFG via project 499180199 (FOR 5522) and via project 492547816 (TRR 360).
@@
The data to generate all figures in this article is available in Zenodo [34].
@@ Appendix A Classical Dynamics of Free Quantum Hard Disks
Appendix B Quantum Many-Body Cages Across the Entire Hilbert Space
Fig. 5 illustrates the entanglement entropy as a function of eigenenergy density, with data points colored by the normalized EA order parameter. This visualization specifically focuses on the largest fragment within the weak fragmentation regime. To extend this analysis, Fig. 7 presents the same plot of entanglement entropy versus eigenenergy density for all fragments. This reveals that numerous additional cages emerge from the smaller fragments, complementing those found in the largest fragment. Notably, the cages observed in the largest fragment primarily stem from quantum interference effects, whereas those in the smaller fragments emerge due to the fragmentation structure itself.
@@ Appendix B Quantum Many-Body Cages Across the Entire Hilbert Space
Figure 7: Quantum Many-Body Cages across the entire Hilbert space. The entanglement entropy of the eigenstates of all fragments is shown for different interaction strengths: (A) $\lambda=0$, representing free hard disks; (B) $\lambda=0.01$, showing weak interactions; (C) $\lambda=0.1$, illustrating moderate interactions; and (D) $\lambda=0.8$, highlighting strong interaction effects.